# On-sky demonstration of an ultra-fast intensity interferometry instrument utilising hybrid single photon counting detectors

### v1

@@ Abstract
Intensity interferometry is a reemerging astronomical technique for performing high angular resolution studies at visible wavelengths, benefiting immensely from the recent improvements in (single) photon detection instrumentation. Contrary to direct imaging or amplitude interferometry, intensity interferometry correlates light intensities rather than light amplitudes, circumventing atmospheric seeing limitations at the cost of reduced sensitivity. In this paper we present measurements with the 1.041.041.04\,1.04m Omicron telescope of C2PU (Centre Pédagogique Planète Univers) at the Calern Observatory in the south of France featuring hybrid single photon counting detectors (HPDs). We successfully measured photon bunching from temporal correlations of three different A-type stars - Vega, Altair and Deneb - in the blue at 405 nm. In all cases the observed coherence time fits well to both the pre-calculated expectations as well as the values measured in preceding laboratory tests. The best signal to noise ratio (SNR), with a value of 12, is obtained for Vega for an observation time of 12.112.112.1\,12.1h. The combination of HPDs and time to digital converter (TDC) results in a timing jitter of the detection system $<50\,$ps. Our setup demonstrates stable and efficient detection of the starlight owed to the large active area of the HPDs. Utilizing a new class of large area single photon detectors based on multichannel plate amplification, high resolution spatial intensity interferometry experiments are within reach at 111\,1m diameter class telescopes within one night of observation time for bright stars.
keywords:
@@
pubyear: 2023pagerange: On-sky demonstration of an ultra-fast intensity interferometry instrument utilising hybrid single photon counting detectorsC
@@ 1 Introduction
The history of stellar intensity interferometry starts with the growing interest in radar technology during the second world war. During that time A.J.F. Siegert looked at the fluctuations in signals finding a relation between the spatio-temporal correlations of the electric fields and of the intensities, today often called the Siegert relation, widely used in intensity interferometry (Siegert, 1943). Using this relation, Hanbury Brown and Twiss proposed a new type of interferometer to determine the diameter of thermal sources such as stars (Hanbury Brown & Twiss, 1954) by measuring the correlation between photons in two separate beams of light (Hanbury Brown & Twiss, 1956). Shortly afterwards they were able to estimate the angular diameter of Sirius A to $0.0063\,^{\prime\prime}$, a unique precision at that time (Hanbury Brown & Twiss, 2013). Later, the first stellar intensity interferometer at the Narrabri Observatory in New South Wales used the described technique in order to measure the angular diameters of 15 stars (Hanbury Brown et al., 1967b; Hanbury Brown et al., 1967a).
@@ 2.1 Stellar Intensity Interferometry
where for the nominator and denominator averages are taken over time (Labeyrie et al., 2006). $I(t)$ denotes the time dependent intensity and $\gamma^{(1)}(\tau)$ is the first order coherence function also referred to as the complex degree of coherence. The second part of Eq. 1 is called the Siegert relation that holds for thermal light sources like stars and connects the second order coherence function to the modulus of the complex degree of coherence (Siegert, 1943). A direct consequence of this relation for thermal light is the effect of "photon bunching". Simply speaking, it is more likely to detect a second photon at a detector when a first photon was already recorded within the light coherence time. For example, for a star $\gamma^{(2)}(\tau\rightarrow\infty)=1$ and therefore $\gamma^{(2)}(\tau=0)>\gamma^{(2)}(\tau\rightarrow\infty)$ holds with $\tau$ the delay in time. The width of this intensity correlation excess around zero time delay is proportional to the star’s coherence time $\tau_{c}$. Moreover, the Wiener-Khinchin theorem relates the electric field spectrum $S(\omega)$ of a time-dependent wave to the Fourier transformation of its temporal autocorrelation function (Labeyrie et al., 2006). Hence, we can express the first order coherence function $\gamma^{(1)}(\tau)$ as follows
@@
This convolution leads to a peak with an amplitude $<1$ and a shape similar to the timing jitter. The area under the curve, and thus the coherence time, are however unchanged if $P_{\mathrm{diff}}(t)$ is properly normalised. Note that, when considering only a single polarisation mode, the height of the bunching peak is reduced to $\sim\tau_{c}/\Delta t$. Using the relation $\tau_{c}=\lambda^{2}/(c\times\Delta\lambda)$ (Mandel & Wolf, 1995), we can estimate the coherence time of any bandpass larger than $\Delta\lambda=0.5\,$nm (FWHM) to $\tau_{c}\ll 1\,$ps. As typical detection systems have a timing jitter around ns to 100100100\,100ps, the detected signal amplitude will be drastically reduced.
Resolving such a signal with diminished amplitude within a correlation histogram can be a major challenge. The minimal amount of observation time necessary to yield a certain signal to noise ratio (SNR) can be estimated under the following assumptions: 1) The observed count rates $n$ are equal for all detectors and constant over the whole measurement time $t_{\mathrm{meas}}$; 2) The bin error for a histogram bin with $N$ clicks is $\Delta N=\sqrt{N}$ (shot noise); 3) The measurement has a timing resolution $\Delta t$ and is binned into a correlation histogram with bin size $t_{\mathrm{bin}}$; 4) the shot noise error is equal for all bins comprising the bunching peak. Propagating the errors we can write the SNR as
$\mathrm{SNR}=\frac{\tau_{c}}{\Delta\tau_{c}}=\frac{\tau_{c}}{t_{\mathrm{bin}}\sqrt{\sum_{i=1}^{m}\left(\Delta\gamma^{(2)}(\tau_{i})\right)^{2}}}.$ (5)
@@
$\mathrm{SNR}=\tau_{c}\frac{1}{t_{\mathrm{bin}}\sqrt{m}}\sqrt{n^{2}\,t_{\mathrm{meas}}\,t_{\mathrm{bin}}}=\tau_{c}\,n\sqrt{\frac{t_{\mathrm{meas}}}{\Delta t}},$ (6)
where we used the fact that the number of bins $m$ making up the correlation peak is equal to the ratio of the timing resolution $\Delta t$ and the bin size $t_{\mathrm{bin}}$ in the limit $\Delta t\gg\tau_{c}$. For this estimation to hold one has to assume the timing resolution to be twice the measured FWHM resolution to account for full sampling of the bunching peak. It should be pointed out that the count rate $n$ is proportional, whereas the coherence time $\tau_{c}$ is inversely proportional, to the optical bandwidth of the observed light field, hence the SNR is independent of the optical bandwidth to first order. When substituting the conventional definition of the coherence time and spectral flux one arrives at the result of (Hanbury Brown & Twiss, 1957).
Finally, Eq. 6 can be used to estimate the necessary measurement time for a given SNR. Assuming a coherence time of 0.20.20.2\,0.2ps (as expected by our setup), a count rate of 222\,2MHz, a detection resolution of 404040\,40ps and aiming for an SNR of 5, we estimate a measurement time of $t_{\mathrm{meas}}\approx 6.8\,$h.
2.2 Omicron telescope
@@ 2.3 Experimental setup
The setup used in order to measure temporal correlations is shown in Fig. 1. The light from the telescope enters the setup from the left and is collimated by an achromatic lens with a focus of $-100\,$mm resulting in a beam diameter of 888\,8mm. A shortpass dichroic mirror (D1) with a cutoff wavelength of 490490490\,490nm transmits the light used for the intensity correlations. The reflected light is focused via an achromatic lens with a focus of 404040\,40mm onto a ZWO ASI120MM-S camera (ZWO) used for guiding. The transmitted light is passed through a polarizing beam splitter (PBS) which is useful in order to increase the contrast of the bunching peak. Besides it separates light for wavefront analysis without a discernible loss in measurement SNR. The light reflected by the PBS impinges on a self-made Shack-Hartmann sensor. A two lens system is used in order to reduce the beam diameter before a $10\times 10\,$mm microlens array (MLA) with a lens pitch of $150\,\mu$m and a focal length of 5.65.65.6\,5.6mm. A second ZWO is placed in the focal plane of the MLA monitoring the spot diagram, i.e. the collimation quality of the observed wavefront. This is an important information as an uncollimated beam would degrade the ultra narrow bandpass. On the main axis the light is first pre-filtered by a Thorlabs bandpass with a central wavelength of 405405405\,405nm and 101010\,10nm bandwidth before passing through a 222\,2nm ultra narrow bandpass with a central wavelength of 405405405\,405nm by Alluxa, both shown as rectangles in Fig. 1. This combination is crucial to minimize the possible out-of band transmission. A $50:50$ non-polarizing beamsplitter (BS) separates the light onto the two hybrid single photon counting detectors (HPDs) from Becker&Hickl with a quantum efficiency of $22.7\,\%$, see measurement plotted in Appendix C. The signal output from the HPDs is fed through a Constant Fraction Discriminator (CFD) produced by Photonscore to generate NIM pulses. The CFD is then connected to a time to digital converter (TDC) from quTools (quTAG). Due to technical limitations the quTAG introduces a dead time of 404040\,40ns much longer than the 555\,5ns dead time of the detectors. With this rather large dead time start-stop correlations are sufficient for histogramming the photon delay times. In order to account for the 666\,6mm active area of the HPDs, lenses with a focal length of 404040\,40mm slowly converge the beam towards the photocathode, shown as red line inside the HPD. Hence, we are not focusing onto the photocathode but rather only gently decrease the beam diameter to fit the size of the cathode leading to a better beam stability. The CFD threshold is set to a similar value as suggested by the HPD manufacturer, whereas the zero crossing is optimized to obtain a high time resolution of around 303030\,30ps. These two settings lead to dark counts on the order of 505050\,50Hz. The NIM output signal of the CFD then travels to the quTAG through a 111111\,11m and 888\,8m signal cable respectively to perform the correlations. The cable delay of 333\,3m is necessary for the signal to lie inside a nearly shot-noise limited environment, see Appendix B. Post-processing is done via a PC.
@@
A CAD render of the setup and the necessary support structure is shown in Fig. 2. The support structure is mounted directly to the output telescope flange, see Fig. 3. The light from the telescope directly enters the setup having a beam diameter of 888\,8mm when passing through the first lens. The HPDs are movable in the xy-plane, using two stepper motors per detector. Note that due to the high stability of the setup the detector positions once aligned and fixed in the lab, stayed in this position throughout the entire measurement. The setup is designed to fit the constraints of the telescope, i.e. weight and dimensions. Counter weights have been used to balance the telescope again after mounting the setup with a weight of approximately 272727\,27kg.
@@
Figure 1: Schematic of the used setup. The light enters from the left. The color splitted yellow light hits the ZWO ASI120MM-S guiding camera and the polarized light is used in the Shack-Hartmann arm in order to monitor the collimation of the wavefront. On the main axis the light passes through two ultra narrow bandpass filters with a central wavelength of 405405405\,405nm before being split onto the two hybrid single photon counting detectors (HPDs). The signal output is fed through a constant fraction discriminator (CFD) and connected to the time to digital converter (TDC). Note that the signal cables have a length difference of 333\,3m.
@@
Figure 2: CAD render of the used setup and its support structure. The light enters through the hole on the left. The ring at the frontplate is used to center the setup inside the telescope flange, see also Fig. 3. One of the aluminium profiles is transparent to enable a better view onto the setup. The optical components are mounted to the groundplate, see Fig. 1 for a detailed explanation of the optical setup. The hybrid single photon counting detectors (HPDs) are motorized via two stepper motors enabling movement in the xy-plane.
@@ 2.4 $2f-2f$ configuration test
Figure 4: Schematic of the $2f-2f$ configuration and the used pupil diaphragm. Light from the xenon short-arc lamp (XBO) is coupled into a single mode fibre (SM) and then into a multi mode fibre (MM) before diverging until the telescope simulator (pupil diaphragm). The resulting mode is similar to the beam profile expected at the Omicron telescope. A lens with a focal length of 200200200\,200mm focuses the light into the experimental setup. In order to tune the count rate, ND-filters can be placed directly after the Source.
Before on-sky observations, the setup was pre-aligned and tested in the lab. An Osram XBO $75$ W$/2$ OFR, i.e. a xenon short-arc lamp (XBO), was used as an artificial star. In order to obtain a similar beam profile as expected at the telescope a $2f-2f$ configuration was used, see Fig. 4. The light from the XBO is first coupled into a single mode fibre in order to collect only a single spatial mode. This single mode fibre is coupled directly to a short piece of $50\,\mu$m mode field diameter step index multimode fibre. The tip of the multimode fibre is used as source from which the light diverges freely. Its diameter is chosen such that it emulates a $25^{\prime}$ field of view, while the short length suppresses the occurrence of adverse timing effects caused by modal dispersion. The diverging beam hits a 3D-printed pupil diaphragm resembling the constraints at the Omicron telescope. The pupil diaphragm, also referred to as the telescope simulator, has a cut out circle with a diameter of $D=17\,$mm and a blocking inner circle with a diameter of 4.94.94.9\,4.9mm to simulate the secondary mirror, see inset of Fig. 4. The resulting mode is focused via a lens with a focal length of 200200200\,200mm. The experimental setup is placed around the focus obtaining a collimated beam after its first lens. Note that all components of the $2f-2f$ configuration were fixed to a 151515\,15mm range micrometer stage. This has the advantage of tuning the focus, hence, facilitating the alignment. The count rate at the HPDs was adjusted via ND-filters that were placed directly after the source.

### v2

# An ultra-fast intensity interferometry instrument

@@ Abstract
Context. Intensity interferometry is a reemerging astronomical technique for performing high angular resolution studies at visible wavelengths, benefiting immensely from the recent improvements in (single) photon detection instrumentation. Contrary to direct imaging or amplitude interferometry, intensity interferometry correlates light intensities rather than light amplitudes, circumventing atmospheric seeing limitations at the cost of reduced sensitivity.
Aims. We aim to develop an ultra-fast, single photon counting and highly stable intensity interferometry instrument for 1 m class optical telescopes. The instrument is supposed to record on sky the expected stellar photon rates and reach the temporal coherence times as measured in the laboratory. In addition, all components, especially the photon detection hardware, of the instrument should be easily upgradeable with custom hardware currently being developed.
Methods. The collimated telescope output is spectrally filtered via an ultra narrow band pass of 2 nm at a central wavelength of 405 nm. We use hybrid photon detectors (HPDs) for single photon detection and a constant fraction discriminator (CFD) for signal conditioning. A time to digital converter (TDC) is used for time stamping. The combination of HPDs and CFDs is optimized for large active area and high timing resolution.
Results. We successfully measured photon bunching of three bright A-type stars - Vega, Altair and Deneb at the 1.041.041.04\,1.04m Omicron telescope of C2PU (Centre Pédagogique Planète Univers) at the Calern Observatory in the south of France. In all cases the observed coherence time fits well to both the pre-calculated expectations as well as the values measured in preceding laboratory tests. We obtained the previously estimated photon count rates at the telescope and achieved highly stable coupling of the star light to the detectors.
Conclusions. Utilizing a new class of large area single photon detectors based on multichannel plate amplification, high resolution spatial intensity interferometry experiments are within reach at 111\,1m diameter class telescopes within one night of observation time for bright stars.
Key Words.:
@@ 1 Introduction
The history of stellar intensity interferometry starts with the growing interest in radar technology during the second world war. During that time A.J.F. Siegert looked at the fluctuations in signals finding a relation between the spatio-temporal correlations of the electric fields and of the intensities, today often called the Siegert relation, widely used in intensity interferometry Siegert (1943). Using this relation, Hanbury Brown and Twiss proposed a new type of interferometer to determine the diameter of thermal sources such as stars (Hanbury Brown & Twiss, 1954) by measuring the correlation between photons in two separate beams of light (Hanbury Brown & Twiss, 1956). Shortly afterwards they were able to estimate the angular diameter of Sirius A to $0.0063\,^{\prime\prime}$, a unique precision at that time (Hanbury Brown & Twiss, 2013). Later, the first stellar intensity interferometer at the Narrabri Observatory in New South Wales used the described technique in order to measure the angular diameters of 15 stars (Hanbury Brown et al., 1967a, b).
@@ 2.1 Stellar Intensity Interferometry
where for the nominator and denominator averages are taken over time (Labeyrie et al., 2006). $I(t)$ denotes the time dependent intensity and $\gamma^{(1)}(\tau)$ is the first order coherence function also referred to as the complex degree of coherence which can be directly measured in amplitude interferometry. The second part of Eq. 1 is called the Siegert relation that holds for thermal light sources such as stars and connects the second order coherence function to the modulus of the complex degree of coherence (Siegert, 1943). A direct consequence of this relation for thermal light is the effect of ”photon bunching”. Simply speaking, it is more likely to detect a second photon at a detector when a first photon was already recorded within the light coherence time. For example, for a star $\gamma^{(2)}(\tau\rightarrow\infty)=1$ and therefore $\gamma^{(2)}(\tau=0)>\gamma^{(2)}(\tau\rightarrow\infty)$ holds with $\tau$ the delay in time. The width of this intensity correlation excess around zero time delay is proportional to the star’s coherence time $\tau_{c}$. Moreover, the Wiener-Khinchin theorem relates the electric field spectrum $S(\omega)$ of a time-dependent wave to the Fourier transformation of its temporal autocorrelation function (Labeyrie et al., 2006). Hence, we can express the first order coherence function $\gamma^{(1)}(\tau)$ as follows
@@
This convolution leads to a peak with an amplitude $<1$ and a shape similar to the timing jitter. The area under the curve, and thus the coherence time, are however unchanged if $P_{\mathrm{diff}}(t)$ is properly normalized. Note that, when considering only a single polarization mode, the height of the bunching peak is reduced to $\sim\tau_{c}/\Delta t$. Using the relation $\tau_{c}=\lambda^{2}/(c\times\Delta\lambda)$ (Mandel & Wolf, 1995), we can estimate the coherence time of any band pass larger than $\Delta\lambda=0.5\,$nm (FWHM) to $\tau_{c}\ll 1\,$ps. As typical detection systems have a timing jitter around ns to 100100100\,100ps, the detected signal amplitude will be drastically reduced. However, our detection system has a much lower timing jitter of 42.442.442.4\,42.4ps resulting in a better timing resolution, when compared with similar measurements.
Resolving such a signal with diminished amplitude within a correlation histogram can be a major challenge. The minimal amount of observation time necessary to yield a certain signal to noise ratio (S/N) can be estimated under the following assumptions: 1) The observed count rates $n$ are equal for all detectors and constant over the whole measurement time $t_{\mathrm{meas}}$; 2) The bin error for a histogram bin with $N$ clicks is $\Delta N=\sqrt{N}$ (shot noise); 3) The measurement has a timing resolution $\Delta t$ and is binned into a correlation histogram with bin size $t_{\mathrm{bin}}$; 4) the shot noise error is equal for all bins comprising the bunching peak. Propagating the errors we can write the S/N as
$\mathrm{S/N}=\frac{\tau_{c}}{\Delta\tau_{c}}=\frac{\tau_{c}}{t_{\mathrm{bin}}\sqrt{\sum_{i=1}^{m}\left(\Delta\gamma^{(2)}(\tau_{i})\right)^{2}}}.$ (5)
@@
$\mathrm{S/N}=\tau_{c}\frac{1}{t_{\mathrm{bin}}\sqrt{m}}\sqrt{n^{2}\,t_{\mathrm{meas}}\,t_{\mathrm{bin}}}=\tau_{c}\,n\sqrt{\frac{t_{\mathrm{meas}}}{\Delta t}},$ (6)
where we used the fact that the number of bins $m$ making up the correlation peak is equal to the ratio of the timing resolution $\Delta t$ and the bin size $t_{\mathrm{bin}}$ in the limit $\Delta t\gg\tau_{c}$. For this estimation to hold one has to assume the timing resolution to be twice the measured FWHM resolution to account for full sampling of the bunching peak. It should be pointed out that the count rate $n$ is proportional, whereas the coherence time $\tau_{c}$ is inversely proportional, to the optical bandwidth of the observed light field, hence the S/N is independent of the optical bandwidth to first order. Nevertheless, the usage of a narrow band pass can be beneficial to increase the height of the correlation peak above systematics inherent to the photon detection system. When substituting the conventional definition of the coherence time and spectral flux one arrives at the result of (Hanbury Brown & Twiss, 1957).
Finally, Eq. 6 can be used to estimate the necessary measurement time for a given S/N. Assuming a coherence time of 0.20.20.2\,0.2ps (as expected by our setup), a count rate of 222\,2MHz, a detection resolution of 404040\,40ps and aiming for an S/N of 5, we estimate a measurement time of $t_{\mathrm{meas}}\approx 6.8\,$h.
2.2 The Omicron telescope
@@ 2.3 Experimental setup
The setup used in order to measure temporal correlations is shown in Fig. 1. The light from the telescope enters the setup from the left and is collimated by an achromatic lens with a focus of $-100\,$mm resulting in a beam diameter of 888\,8mm. A short pass dichroic mirror (D1) with a cutoff wavelength of 490490490\,490nm transmits the light used for the intensity correlations. The reflected light is focused via an achromatic lens with a focus of 404040\,40mm onto a ZWO ASI120MM-S camera (ZWO) used for guiding. The transmitted light is passed through a polarizing beam splitter (PBS) which is useful in order to increase the contrast of the bunching peak. Besides it separates light for wavefront analysis without a discernible loss in measurement S/N. The light reflected by the PBS impinges on a self-made Shack-Hartmann sensor. A two lens system is used in order to reduce the beam diameter before a $10\times 10\,$mm microlens array (MLA) with a lens pitch of $150\,\mu$m and a focal length of 5.65.65.6\,5.6mm. A second ZWO is placed in the focal plane of the MLA monitoring the spot diagram, that is the collimation quality of the observed wavefront. This is an important information as an uncollimated beam would degrade the ultra narrow band pass. On the main axis the light is first pre-filtered by a Thorlabs band pass with a central wavelength of 405405405\,405nm and 101010\,10nm bandwidth before passing through a 222\,2nm ultra narrow band pass with a central wavelength of 405405405\,405nm by Alluxa, both shown as rectangles in Fig. 1. This combination is crucial to minimize the possible out-of band transmission. A $50:50$ non-polarizing beam splitter (BS) separates the light onto the two hybrid single photon counting detectors (HPDs) from Becker&Hickl with a quantum efficiency of $22.7\,\%$, see measurement plotted in Appendix C. The signal output from the HPDs is fed through a Constant Fraction Discriminator (CFD) produced by Photonscore to generate NIM pulses. The CFD is then connected to a time to digital converter (TDC) from quTools (quTAG). Due to technical limitations the quTAG introduces a dead time of 404040\,40ns much longer than the 555\,5ns dead time of the detectors. With this rather large dead time start-stop correlations are sufficient for histogramming the photon delay times. In order to account for the 666\,6mm active area of the HPDs, lenses with a focal length of 404040\,40mm slowly converge the beam towards the photocathode, shown as red line inside the HPD. Hence, we are not focusing onto the photocathode but rather only gently decrease the beam diameter to fit the size of the cathode leading to a better beam stability. The CFD threshold is set to a similar value as suggested by the HPD manufacturer, whereas the zero crossing is optimized to obtain a high time resolution of around 303030\,30ps. These two settings lead to dark counts on the order of 505050\,50Hz. The NIM output signal of the CFD then travels to the quTAG through a 111111\,11m and 888\,8m signal cable respectively to perform the correlations. The cable delay of 333\,3m is necessary for the signal to lie inside a nearly shot-noise limited environment, see Appendix B. Post-processing is done via a PC.
@@
A CAD render of the setup and the necessary support structure is shown in Fig. 2. The support structure is mounted directly to the output telescope flange, see Fig. 3. The light from the telescope directly enters the setup having a beam diameter of 888\,8mm when passing through the first lens. The HPDs are movable in the xy-plane, using two stepper motors per detector. Note that due to the high stability of the setup the detector positions once aligned and fixed in the lab, stayed in this position throughout the entire measurement. The setup is designed to fit the constraints of the telescope, these are. weight and dimensions. Counter weights have been used to balance the telescope again after mounting the setup with a weight of approximately 272727\,27kg.
@@
Figure 1: Schematic of the used setup. The light enters from the left. The color split yellow light hits the ZWO ASI120MM-S guiding camera and the polarized light is used in the Shack-Hartmann arm in order to monitor the collimation of the wavefront. On the main axis the light passes through two ultra narrow band pass filters with a central wavelength of 405405405\,405nm before being split onto the two hybrid single photon counting detectors (HPDs). The signal output is fed through a constant fraction discriminator (CFD) and connected to the time to digital converter (TDC). Note that the signal cables have a length difference of 333\,3m.
@@
Figure 2: CAD render of the used setup and its support structure. The light enters through the hole on the left. The ring at the front plate is used to center the setup inside the telescope flange, see also Fig. 3. One of the aluminum profiles is transparent to enable a better view onto the setup. The optical components are mounted to the ground plate, see Fig. 1 for a detailed explanation of the optical setup. The hybrid single photon counting detectors (HPDs) are motorized via two stepper motors enabling movement in the xy-plane.
@@ 2.4 $2f-2f$ configuration test
Figure 4: Schematic of the $2f-2f$ configuration and the used pupil diaphragm. Light from the xenon short-arc lamp (XBO) is coupled into a single mode fiber (SM) and then into a multi mode fiber (MM) before diverging until the telescope simulator (pupil diaphragm). The resulting mode is similar to the beam profile expected at the Omicron telescope. A lens with a focal length of 200200200\,200mm focuses the light into the experimental setup. In order to tune the count rate, ND-filters can be placed directly after the Source.
Before on-sky observations, the setup was pre-aligned and tested in the lab. An Osram XBO $75$ W$/2$ OFR, that is a xenon short-arc lamp (XBO), was used as an artificial star. In order to obtain a similar beam profile as expected at the telescope a $2f-2f$ configuration was used, see Fig. 4. The light from the XBO is first coupled into a single mode fiber in order to collect only a single spatial mode. This single mode fiber is coupled directly to a short piece of $50\,\mu$m mode field diameter step index multimode fiber. The tip of the multimode fiber is used as source from which the light diverges freely. Its diameter is chosen such that it emulates a $25^{\prime}$ field of view, while the short length suppresses the occurrence of adverse timing effects caused by modal dispersion. The diverging beam hits a 3D-printed pupil diaphragm resembling the constraints at the Omicron telescope. The pupil diaphragm, also referred to as the telescope simulator, has a cut out circle with a diameter of $D=17\,$mm and a blocking inner circle with a diameter of 4.94.94.9\,4.9mm to simulate the secondary mirror, see inset of Fig. 4. The resulting mode is focused via a lens with a focal length of 200200200\,200mm. The experimental setup is placed around the focus obtaining a collimated beam after its first lens. Note that all components of the $2f-2f$ configuration were fixed to a 151515\,15mm range micrometer stage. This has the advantage of tuning the focus, hence, facilitating the alignment. The count rate at the HPDs was adjusted via ND-filters that were placed directly after the source.